### v4

# SCMAlgebras: a Macaulay2 package to check sequential Cohen-Macaulayness

Abstract.
We introduce the Macaulay2 package SCMAlgebras. It provides functions for computing the modules of deficiency and the filter ideals, in order to check whether a module or an ideal is sequentially Cohen-Macaulay. After the basic algebraic notions and results are recalled, the main features of the package are described through examples.
Key words and phrases:
sequentially Cohen-Macaulay, modules of deficiency, filter ideals, unmixed layers, unmixed
1991 Mathematics Subject Classification:
Primary 13H10 Secondary 13D07, 13A02, 13F20
SCMAlgebras v1.1
1. Introduction
The notion of sequentially Cohen-Macaulay modules was originally introduced by Stanley [21] in the graded setting, as an algebraic counterpart to the concept of nonpure shellable complexes, which had been developed by Björner and Wachs [3] in the context of combinatorics and discrete geometry. Later on, Schenzel [19] independently introduced the notion of Cohen-Macaulay filtered modules and proved the equivalence of this notion to the sequentially Cohen-Macaulay property. Since then, the theory of sequentially Cohen-Macaulay modules has garnered significant attention from many authors [1, 2, 5, 6, 7, 8, 9, 11, 13, 15, 20, 22, 23], particularly due to its deep connections with the study of algebraic properties of combinatorial objects such as simplicial complexes and graphs.
In this paper, we present a new Macaulay2 [10] package, SCMAlgebras [14], designed to check whether a given module or ideal satisfies the sequentially Cohen-Macaulay property. This package implements key theoretical results, such as Theorem 2.6, which characterizes sequentially Cohen-Macaulay modules in terms of their modules of deficiency, and Proposition 2.9 due to Goodarzi [9], which provides an efficient criterion for verifying the sequentially Cohen-Macaulay property for homogeneous ideal. To support these results, the package introduces new computational tools, such as functions for calculating the modules of deficiency (see Definition 2.3) and generating filter ideals (see Definition 2.8), both of which are crucial for testing the sequentially Cohen-Macaulay property of the input objects. Moreover, the filter ideals serve as useful tools for checking the unmixedness of an ideal (Proposition 2.11), via the notion of unmixed layer.
These tools extend the functionality of the software, playing as a starting point for future development in the manipulation of sequentially Cohen-Macaulay property in both theoretical and applied contexts.
2. Mathematical background
In this section we recall the notion of sequentially Cohen-Macaulay modules and the main characterizations, that are implemented in the package. In the following we will work in the graded case. Let $S=K[x_{1},\ldots,x_{n}]$ be the standard graded polynomial ring in $n$ variables over a field $K$.
Definition 2.1.
A finitely generated graded $S$-module $M$ is called a sequentially Cohen-Macaulay module if there exist a finite filtration of submodules of $M$
$0=M_{0}\subset M_{1}\subset\cdots\subset M_{d-1}\subset M_{r}=M$
such that
  1. a)
    each quotient $M_{i}/M_{i-1}$ is Cohen-Macaulay,
  2. b)
    $\dim M_{i}/M_{i-1}<\dim M_{i+1}/M_{i}$ for all $i$.
A filtration satisfying the above conditions is called a sequentially Cohen-Macaulay filtration of $M$. It is easy to see that if the sequentially Cohen-Macaulay filtration exists, then it is unique.
Remark 2.2.
Any Cohen-Macaulay module is sequentially Cohen-Macaulay, by considering the sequentially Cohen-Macaulay filtration $0=M_{0}\subset M_{1}=M$. Hence, one has the implication
{tblr}
colspec=ccc,rows=m,colsep=.4cm,width=rowsep=.5cm Cohen-Macaulay & $\Longrightarrow$ sequentially
Cohen-Macaulay
but the converse is not true in general.
Proving the sequentially Cohen-Macaulay property through the construction of a sequentially Cohen-Macaulay filtration is not always an easy task. The definition itself is not operative and it is not very efficient when trying to implement it on algorithms. In order to demonstrate this condition effectively and one must rely on algebraic or combinatorial characterization.
In the ladder, two of the main characterizations of the sequentially Cohen-Macaulay property are recalled. The first one involves the notion of the modules of deficiency.
Definition 2.3.
Let $M$ be a finitely generated graded $S$-module and let $d=\dim M$. For any $i\in\mathbb{Z}$ the $i$th module of deficiency of $M$ is
$\omega^{i}(M)=\operatorname{Ext}^{n-i}_{S}(M,S(-n)).$
If $i=d$, the module $\omega^{d}(M)=\omega(M)$ is the canonical module of $M$.
Remark 2.4.
The modules of deficiency were introduced and studied by Schenzel in [17] in the local case. By duality (in the graded setting), one has that $\omega^{i}(M)$ is the Matlis dual of the $i$th local cohomology module $H^{i}(M)$ of $M$. Let $t=\operatorname{depth}M$.
From a well–known result of Grothendieck (see [4]), it holds that $H^{i}(M)=0$ for $i<t$ or $i>d$. Moreover, $H^{i}(M)\neq 0$ for $i=t,d$. It follows that $M$ is Cohen-Macaulay if and only if $H^{i}(M)=0$ for all $i\neq d$. Hence, one has $\omega^{i}(M)=0$ for $i<t$ or $i>d$, $\omega^{t}(M)\neq 0$ and $\omega(M)=\omega(M)\neq 0$ and $M$ is a Cohen-Macaulay module if and only if $\omega^{i}(M)=0$ for all $i\neq d$. In such case, the canonical module $\omega^{d}(M)=\omega(M)$ is also Cohen-Macaulay. Hence, the modules of deficiency of $M$ measure how far is $M$ from being a Cohen-Macaulay module.
Definition 2.5.
A module $M$ with Cohen-Macaulay canonical module $\omega(M)$ is called a canonically Cohen-Macaulay module.
The next classical result characterizes sequentially Cohen-Macaulay modules via the modules of deficiency (see [6]).
Theorem 2.6.
Let $M$ be a finitely generated $S$-module, with $d=\dim M$. Then, the followings are equivalent:
  1. 1.
    $M$ is sequentially Cohen-Macaulay;
  2. 2.
    for all $0\leq i\leq d$ the $i$th module of deficiency $\omega^{i}(M)$ is either zero or an $i$-dimensional Cohen-Macaulay module;
  3. 3.
    for all $0\leq i<d$ the $i$th module of deficiency $\omega^{i}(M)$ is either zero or an $i$-dimensional Cohen-Macaulay module.
The equivalence 1. $\Longleftrightarrow$ 2. in the above theorem was announced without proof in [21, Theorem 2.11], but the author mentioned a spectral sequence argument due to Peskine.
Remark 2.7.
If $M$ is a sequentially Cohen-Macaulay module, then by condition 2. of Theorem 2.6 $M$ has a Cohen-Macaulay canonical module. Hence, it follows that a sequentially Cohen-Macaulay module is canonically Cohen-Macaulay, but the converse is not true in general.
Thus, merging the informations from Remark 2.2 and Remark 2.7, one has the following chain of implications:
{tblr}
colspec=ccccc,rows=m,colsep=.4cm,width=rowsep=.5cm Cohen-Macaulay & $\Longrightarrow$ sequentially
Cohen-Macaulay $\Longrightarrow$ canonically
Cohen-Macaulay
with the reverse implications not true in general.
Although Theorem 2.6 remains useful in many situations (see [15, Theorem 1.5],[20]), it also presents two significant challenges. Firstly, to determine which modules of deficiency vanish is often a difficult task (see [23]). Secondly, proving the Cohen-Macaulayness of the non–vanishing modules is typically non–trivial. These difficulties underscore the complexity of applying the theorem in practice, particularly when dealing with more complicated algebraic structures.
In case we work with a homogeneous ideal $I\subset S$ it is possible to characterize the sequentially Cohen-Macaulay property by using new ideals, namely the filter ideals, which are constructed from the primary decomposition of $I$.
Definition 2.8.
Let $I\subset S$ be a homogeneous ideal, with $d=\dim S/I$, and let $I=\bigcap_{j=1}^{r}Q_{j}$ be the minimal primary decomposition of $I$. For all $1\leq j\leq r$, let $P_{j}=\sqrt{Q_{j}}$ be the radical of $Q_{j}$. For all $-1\leq i\leq d$, the $i$th filter ideal of $I$ is
$I^{<i>}=\bigcap_{\dim S/{P_{j}}>i}Q_{j},$
where $I^{<-1>}=I$ and $I^{<d>}=S$.
Moreover, we define the minimum dimension of $I$ as the minimum integer such that $I^{<i>}\neq I$. By definition, it is clear that minimum dimension is at least $-1$.
For a homogeneous ideal $I\subset S$, one has the filtration
$I=I^{<-1>}\subseteq I^{<0>}\subseteq I^{<1>}\subseteq\ldots\subseteq I^{<d-1>}\subseteq I^{<d>}=S.$
Moreover, rewriting the above sequence of inclusions modulo $I$, we obtain
$0\subseteq I^{<0>}/I\subseteq I^{<1>}/I\subseteq\ldots\subseteq I^{<d-1>}/I\subseteq S/I,$
which is called the dimension filtration of $S/I$. One can show (see [9, Lemma 1]) that such filtration does not depend on the primary components, therefore it is unique. Note that, the filter ideals $I^{<i>}$ are also independent from the primary decomposition.
By the above construction, one has the following characterization of the sequentially Cohen-Macaulay property for $S/I$ due to Goodarzi (see [9, Proposition 16]).
Proposition 2.9.
Let $I\subset S$ be a homogeneous ideal and suppose $d=\dim S/I$. Then, $S/I$ is sequentially Cohen-Macaulay if and only if
$\operatorname{depth}S/{I^{<i>}}\geq i+1,\qquad\text{for all $0\leq i<d$}.$ (1)
Goodarzi’s characterization provides an efficient criterion to verify the sequentially Cohen-Macaulay property for a homogeneous ideal. It has been widely used, for instance, to study the sequential Cohen-Macaulayness of the binomial edge ideals of various classes of graphs, such as closed graphs (see [8]), block graphs, cycles, and cone graphs (see [15]), playing a crucial role in the proofs. In particular, for binomial edge ideals, this result also permitted in [15] the investigation of situations that were difficult to describe using Theorem 2.6. This breakthrough allowed for a deeper understanding of cases previously out of reach, thus highlighting the power of Goodarzi’s approach.
We end this section with another useful application of filter ideals.
Definition 2.10.
et $I\subset S$ be a homogeneous ideal, with $d=\dim S/I$. For all $1\leq i\leq d$, the $i$th unmixed layer of $S/I$ is the $S$-module $I^{<i>}/I^{<i-1>}$ and it is denoted by $U_{i}(S/I)$.
The name unmixed layer is motivated by [19, Corollary 2.3], in fact each module $U_{i}(S/I)$ is unmixed, that is for all $i$ the associated primes of $U_{i}(S/I)$ have the same height. The next result (see [9, Lemma 3]) gives a characterization of the unmixedness of the ideal $I$.
Proposition 2.11.
Let $I\subset S$ be a homogeneous ideal, with $d=\dim S/I$. Then $S/I$ is unmixed if and only if $U_{i}(S/I)=0$ for all $1\leq i<d$ and $U_{d}(S/I)=S/I$.
3. The package
In this section we describe the functions of the package SCMAlgebras [14] by showing some examples of usage and algorithms.
Example 3.1.
Let $S=K[x_{1},x_{2},x_{3},x_{4},x_{5}]$ and let $M$ be the $S$-module defined as the $\operatorname{Coker}$ of the map represented by the matrix $\left(\begin{smallmatrix}x_{1}x_{2}&x_{3}x_{4}&0&0\\ 0&x_{1}x_{5}&x_{2}x_{4}&0\end{smallmatrix}\right)$. Such a module has $\operatorname{depth}M=3$ and $\dim M=4$. We use the deficiencyModule function to compute the modules of deficiency of $M$.
i1 : loadPackage "SCMAlgebras";
i2 : S=QQ[x_1..x_5];
i3 : M=coker matrix{{x_1*x_2,x_3*x_4,0,0},{0,x_1*x_5,x_2*x_4,0}}
o3 = cokernel | x x x x 0 0 |
1 2 3 4
| 0 x x x x 0 |
1 5 2 4
2
o3 : S-module, quotient of S
i4 : deficiencyModule(M,1)
o4 = 0
o4 : S-module
i5 : deficiencyModule(M,3)
2 2
o5 = cokernel | x x x x x x x |
1 5 3 4 1 2 4
1
o5 : S-module, quotient of S
The function deficiencyModule(M,i) uses Definition 2.3 and computes the module $\operatorname{Ext}^{n-i}_{S}(M,S(-n))$. According to Remark 2.4, in our case, the function returns the zero module whenever $i<3$ or $i>4$, and the only non vanishing modules of deficiency are $\omega^{3}(M)$ and $\omega^{4}(M)=\omega(M)$, i.e. the canonical module of $M$. Moreover, the function canonicalModule(M) allows to compute the canonical module $\omega(M)$ directly, as the module of deficiency corresponding to dimension.
i6 : canonicalModule(M) == deficiencyModule(M,4)
o6 = true
Hence, one can check if $M$ is sequentially Cohen-Macaulay using the isSCM function, involving the modules of deficiency, as seen in Theorem 2.6.
i7 : isSCM M
o7 = false
Moreover, using the isCCM function, one can check that $M$ is canonically Cohen-Macaulay, by checking the Cohen-Macaulayness of its canonical module.
i8 : isCCM M
o8 = true
Remark 3.2.
Note that, while in Example 3.1, the functions deficiencyModule and canonicalModule are applied to a module, they can also be applied to an ideal.
Example 3.3.
Consider the graph $G$ depicted below.
[…]
Consider the ideal $J_{G}$ in the polynomial ring $S=K[x_{1},\ldots,x_{10},y_{1},\ldots,y_{10}]$ generated by all the binomials $f_{ij}=x_{i}y_{j}-x_{j}y_{i}$, such that $\{i,j\}$ with $i<j$ is an edge of $G$, namely the binomial edge ideal (see [12, 16]) of $G$.
i1 : loadPackage "SCMAlgebras";
i2 : S = QQ[x_1..x_10,y_1..y_10];
i3 : E = {{1,2},{1,3},{1,4},{1,5},{1,6},{1,7},{1,8},{1,9},{1,10},
{6,7},{8,9},{8,10},{9,10}};
i4 : J=ideal(for e in E list x_(e#0)*y_(e#1)-x_(e#1)*y_(e#0));
The function filterIdeals computes the filter ideals of $J_{G}$, according to Definition 2.8. Moreover, the function minimumDimension returns the smallest integer such that the filter ideals differ from $J_{G}$.
i5 : filterIdeal(J,13)
o5 = ideal (y , x , x y - x y , x y - x y , x y - x y ,
1 1 10 9 9 10 10 8 8 10 9 8 8 9
x y - x y )
7 6 6 7
o5 : Ideal of S
i6 : minimumDimension J
o6 = 11
i7 : filterIdeal(J,7) == J
o7 = true
i8 : filterIdeal(J,11) == J
o8 = false
Using the function unmixedLayer, one can compute the unmixed layers of $J_{G}$, according to Definition 2.10.
i9 : unmixedLayer(J,13)
o9 = subquotient (| y x x y - x y x y - x y
1 1 10 9 9 10 10 8 8 10
x y - x y x y - x y |, | y x x y - x y
9 8 8 9 7 6 6 7 1 1 10 9 9 10
x y - x y x y - x y x y - x y |)
10 8 8 10 9 8 8 9 7 6 6 7
1
o9 : S-module, quotient of S
To check the unmixedness of $J_{G}$, the function isUnmixed can be applied to the ideal J. In this case Proposition 2.11 is used.
i10 : isUnmixed J
o10 = false
Finally, one can check that $S/J_{G}$ is sequentially Cohen-Macaulay (as proved in [15, Theorem 1.5]) using the isSCM function. This time we apply the function to the ideal J, in such case Proposition 2.9 is used.
i11 : isSCM J
o11 : true
Remark 3.4.
At the end of Example 3.1 we applied the isSCM function to J and Goodarzi’s characterization (Proposition 2.9) was used to check the sequentially Cohen-Macaulay property for $S/J_{G}$. Note that, to make the check involving Theorem 2.6, one can apply the function to S^1/J instead of J,
i10 : isSCM(S^1/J)
o10 : true
getting, as expected, the same result.
We conclude by listing all the functions available in the package SCMAlgebras.
[…]
columns=l,rows=.75cm,m,hline1,2,Z=1.2pt,row1=lightgray {tblr}Q[1] Q[2] Function & Description
deficiencyModule(M,i) Computes the $i$th module of deficiency $\omega^{i}(M)$ of $M$
canonicalModule(M) Computes the canonical module $\omega(M)$ of $M$
minimumDimension(I) Computes the minimum dimension of $I$
filterIdeal(I,i) Computes the $i$th filter ideal $I^{<i>}$ of $I$
unmixedLayer(I,i) Computes the $i$th unmixed layer $U_{i}(I)$ of $I$
isUnmixed(I) Checks if $I$ is unmixed
isSCM(M) Checks if $M$ is sequentially Cohen-Macaulay
isCCM(M) Checks if $M$ is canonically Cohen-Macaulay
Table 1. List of functions of SCMAlgebras v1.1.
4. Conclusion and perspective
The examples and the algorithms presented in this work are part of the Macaulay2 [10] package SCMAlgebras v1.1 [14]. The tests were performed with version 1.25.05 of Macaulay2.
We are confident that the methods and functionalities provided herein may be valuable for further applications. Moreover, the structure of the package allows for future developments, specifically by implementing additional key results and techniques. This task is currently under study by the author.
Acknowledgments
The author thanks Prof. Giancarlo Rinaldo for sharing his original source code, which was later integrated into this package and for his helpful suggestions. The author also thanks Prof. Marilena Crupi her helpful suggestions. The author acknowledge support of the GNSAGA group of INdAM (Italy).